# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ A.2 Financial Institutions in the Cities
Notes: This map shows the location of the pawnshops and moneylenders in Tsukishima. The bullseye indicates the pawnshop (shichiya). The black circle shows the money lenders (kanekashi gyō). The white circle shows the rickshaw station (kuruma yado).
Source: Department of Health, Ministry of the Interior 1923c, tenth map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
@@ A.3 Development of Tsukishima
Figure A.3: Tsukishima: Ishikawajima, Tsukudajima, Shin-tsukudajima, and Lots 1–3
@@
Notes: Tsukishima contains Ishikawajima, Tsukudajima, Shin-tsukudajima, and Tsukishima Lots 1–3. This map does not illustrate the northernmost part of Ishikawa jima (see Figure 3 for this part). “Ferries” shows the Tsukishima to the mainland of Kyobashi Ward sea routes called Tsukuda, Tsukishima, and Kachidoki routes (right to left). “Aioi bridge” links Shin tsukudajima and Fukagawa ward.
Source: Department of Health, Ministry of the Interior 1923c, 15th map. The author modified and adjusted the tone using Adobe Photoshop 24.7.0.
Tsukishima (transliterated as “moon island”) is a reclaimed land on a sandbar of the Sumida River. Ishikawajima and Tsukudajima islands at the northern end of Tsukishima (Figure A.3) are the starting point of its history. The following description in this section is based on Chūō Ward (1994, pp. 122–124).
@@
Ishikawajima was built in 1626 by a boatman, Hachizaemon Ishikawa, who received a fief from the Shogunate. In 1790, the Ishikawajima Prison (ninsoku yoseba) was built on Ishikawajima. This was a facility for the reintegration of minor offenders and those without lodgings. After Matthew Perry arrived in Japan in 1853, the Shogunate ordered the Mito Clan to build a shipyard on Ishikawajima. A month later, a groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856. In 1866, the Ishikawajima Beacon was built on the west bank of the Ishikawajima Prison and served to signal the position of ships entering the Nihonbashi direction. In 1870, the Ishikawajima Prison became a place of imprisonment (toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877. The prison was finally moved to Sugamo in Kitatoshima County in Tokyo prefecture, and the penitentiary on Ishikawajima was closed in 1895.
@@
The area adjacent to Ishikawajima is called Tsukudajima, where fishermen from “Tsukuda” in Osaka prefecture settled in 1644 after receiving land from the Shogunate. In 1645, the “Tsukuda ferry” began, and throughout the Edo period, Tsukudajima was known as a fishing island for whitefish for the Shogun’s household. In 1863, the Tsukudajima Battery was built on the south side of Tsukudajima—one of $13$ batteries used to defend Edo Bay. This battery later became the starting point for the reclamation of Lot 1 of Tsukishima.
@@
In 1884, the Port of Tokyo Waterway Dredging Project was initiated. This was to secure the passage of ships by sweeping away the sediment in the shipping channel, and the Tsukishima landfill was created using the deposit. The first Tsukishima site was completed in 1891, and the “Tsukishima ferry” began the same year. The second site was created in 1894; the Shin Tsukudajima (namely, new Tsukudajima) was built in 1896. From the west shore of Shin Tsukudajima, the island overlooks the Boso Peninsula in Chiba Prefecture, making it a favorite vacation spot for writers. Cultural figures such as Toson Shimazaki and Kaoru Osanai stayed here from 1907 to the early Taisho period. In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to development of industry. [^84] Until then, a retailer named Suzukiiya had been selling water to supply the entire island using a ferryboat. The “Kachidoki Ferry” began to deliver to Lot 2 in 1905. Lot 3 was finally completed in 1913.
@@ A.4 Housing in Tsukishima
The Tsukishima Survey included a housing survey (Table B.1). Table A.4 summarizes the housing types measured in this survey. As shown, $88$ ($46$%) of the detached homes are in Tsukudajima. Tsukudajima is reported to be “superior to other areas in terms of housing quality, and the percentage of inferior housing is meager.” [^85] The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp. 69–76). Shin-tsukudajima had a low percentage of detached homes. The report states that it was indeed “inhabited by unskilled laborers” and “occupied the most inconvenient housing status of all the islands.” Lot 1 was the center of life in Tsukishima and was a cluster of numerous residences, mainly tenements. Lot 2 had a relatively high percentage of single-family and two-family houses and appears to be characterized as a “workers’ residential area” than Lot 1. These indicate that there might have been some geographic differences in income distribution regarding housing type. Overall, however, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) were row houses, and only $190$ were single-family homes. This confirms that Tsukishima was an industrial area with many working families. The workers lived in similar tenements and with similar standards of living. Figure A.4 is a photo showing these tenement houses.
@@
Note: Figures A.4a and A.4b show the two- and one-story tenements in Tsukishima, respectively.
Source: Department of Health, Ministry of the Interior 1923c, figures 62–63; 66.
@@ A.5 Machinery Factories in Tsukishima
First, I used the results of the Primary School Children Survey included in the Tsukishima Survey to analyze the occupation (industry) of the household heads in Tsukishima. This is a survey of 4th-6th graders at Tsukishima 1st and 2nd Elementary Schools in July 1919. There were $1,173$ children in these grades, and the survey included all $1,100$ children who resided in Tsukishima. At that time, there were no other elementary schools. Thus, this survey is a complete survey for all children in grades 4–6 in Tsukishima. Although it does not include households without children in grades 4–6, it is undoubtedly a beneficial survey that provides an overview of the households with children—the main target of this study.
Panel A of Table A.5 shows the industrial structures. The composition ratios are similar to the census-based statistics reported in Panel A of Table 1. It does not specify whether brothers and sisters in the same household were counted twice. Despite this, the similarity confirms that the statistics obtained from this survey are useful to identify overall trends in family households in Tsukishima.
Panel B of Table A.5 divides the “social classes” of the household heads working in the manufacturing sector into six categories. In the report, large business owners were classified into the capitalist class; small business owners, free enterprise owners, and directors were regarded as the middle class; workers were defined as the working class (Department of Health, Ministry of the Interior 1923a, p. 50). Most heads of households were workers, with only 1% of the capitalist class and less than 20% of the middle class. While the definition of “executive” is unclear, “small business owners” must be small factory owners who could have had similar earnings to the skilled worker households (Department of Health, Ministry of the Interior 1923a, p. 64).
@@
Notes: This table summarizes the statistics from the complete survey on the 4th–6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Panel A summarizes the heads’ industries, and Panel B sorts the “worker type” of the heads in the manufacturing sector, respectively.
Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
@@
In November 1920, there were $214$ factories in Tsukishima, $168$ of them (approximately $80$%) were classified as machinery factories. The official report documented that the wooden pattern factories could be included in the machinery sector because they were incorporated into the casting processes (Department of Health, Ministry of the Interior 1923a, p. 388). The share of machinery factories becomes approximately $90$% if I include those $24$ wooden pattern factories. Panel A of Table A.6 uses the manufacturing survey in the Tsukishima Survey to classify the $168$ machinery factories by size and industrial sector. First, more than $90$% of the plants are in the three sectors of machinery and equipment manufacturing, can manufacturing, and blacksmithing, with only $15$ plants in the casting sector. $113$ ($67$%) of the machinery factories have less than $15$ workers. Second, many smithing factories have fewer than $15$ workers. Hence, almost all of Tsukishima’s smithing industry was classified as small businesses. [^86] The report suggests that $43$ of the $47$ smithing factories were “town factories” or “workplaces.” The definition of “town factory” is not clear in the report (p. 392). Since they were always counted more than the number of factories with 15 or fewer workers, they are likely to include smaller factories operated as small businesses.
@@
Panel B of Table A.6 classifies the types of the $620$ workers based on the results of the primary school survey used in Table A.5. There are $455$ skilled and $165$ unskilled workers. The same report shows that $490$ out of $620$ workers are industrial workers. The day laborers are classified as non-industrial workers. In fact, $620$ workers minus $120$ day laborers become $490$, which is identical to the number of industrial workers documented. Therefore, approximately $90$% of factory workers ($455/490$) can be regarded as skilled workers in the industrial sector: $65$% work in factories, $17$% in cottage industries, and $18$% are craftsmen. Since workers in cottage and handicraft industries are classified as working in small factories, $65$% of these workers might have worked in the middle- to large-scale factories (with $15+$ workers).
@@
Notably, this survey only includes household heads with primary school children aged 9–11 (i.e., 4th to 6th grades), suggesting that most of these heads are in their thirties, in the skilled workers age range. Therefore, the proportion of skilled workers to the entire worker group could be smaller than the $90$%. Nevertheless, this confirms that most of the Tsukishima workers in their 30s were skilled workers at that time.
@@
Notes: Panel A summarizes the number of machinery factories by the scale and type of the factory. Statistics are from the manufacturing census conducted in the Tsukishima Survey (November 1, 1920). The threshold of the scale (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916.
Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th–6th graders in Tsukishima1st and 2nd Primary Schools in July 1919.
Sources: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
@@ A.6 Works in the Machinery Factories
Tsukishima Survey includes the factory survey that counts the number of machinery factories in Tsukishima. Panel A of Table A.6 summarizes the number of machinery factories by type of plant. The machinery and equipment manufacturing factories shared $39$% of all the machinery factories. Typical jobs for workers in machinery and equipment manufacturing plants are finishers (or fitters) and turners. A finisher is “a skilled worker responsible for finishing, assembling, and commissioning machine components” and is “one of the finest skilled workers because it requires skills in all aspects to correct defects in machine making” (Department of Health, Ministry of the Interior 1923a, p. 393). The turner is the “middle axis of machinists,” to which most semi-skilled and skilled workers belong. Becoming a full-fledged finisher and lathe operator takes at least five years. Generally, apprentices enter the workforce at approximately 14–18, but they do not work in the same factory for many years within the apprenticeship system. After learning the skills for 1–2 years, they often move to another factory to improve their skills (Department of Health, Ministry of the Interior 1923a, p. 394). Wages are usually paid on a piece-work basis or on a time basis. In Tsukishima, a time contract is broadly adopted, and payment is based on the “residual time” method. The “residual time” method means that “if the work is completed within the contracted time, the remaining time is considered a bonus.” For example, if $12$ hours are contracted and the work is completed in $10$ hours, $2$ hours are paid as a bonus (pp. 396-397). Wages are paid by dividing one month into two periods (the first and second half of the year). The average work hours are reported to be approximately ten hours, with occasional all-night work (p. 402). Overtime hours are approximately two hours per day (Department of Health, Ministry of the Interior 1923a, p. 402).
@@
Next, I summarize labor in the canning factory, which accounts for $24$% of all the factories. The canning factory is engaged in shipbuilding, bridges, steel frames for construction, and other works, including fabrication and repair of gas reservoirs (p. 406). Workers in a canning factory are called “boiler makers” and are typically involved in bending steel plates and jointing them with rivets. Workers are divided into rivetter, holder-up, plater, angle-ironsmith, and caulking. The work is usually done in groups, with the skilled angle-ironsmith and caulking. The “bōsin” of the rivetter gives the instructions as the head of the group. Wages are paid on a piece-rate contract basis. Because of the contracted wage, the cannery workers’ income is unstable; thus, they are paid a fixed wage (daily wage amount) during the off-season.
@@
Smithing is the second most common industry, accounting for $28$% of the total. Panel A of Table A.6 reveals that most of the smithing industry is relatively small in scale. They make boulders, nuts, rivets, lathe stops, jigs, and spanners for can manufacturing and construction. They are subcontractors who manufacture metal parts for the larger factories in Tsukishima (Department of Health, Ministry of the Interior 1923a, p. 412). There are three types of jobs: yokoza, sakite, and tatara-fuki (bellows-blower). The yokoza is a self-employed skilled worker (called oyakata), while the other two are young unskilled laborers. In small operations, wages are based on a percentage based on voluntary business discussions with the proprietor, and working hours are not regular (pp. 411–413).

### v14

@@ A.2 Financial Institutions in the Cities
Notes: This map shows the location of the pawnshops and moneylenders in Tsukishima. The bullseye indicates the pawnshop (shichiya). The black circle shows the money lenders (kanekashi gyō). The white circle shows the rickshaw station (kuruma yado). The upper right corner indicates north. Source: Department of Health, Ministry of the Interior 1923c, tenth map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Moneylenders were another lending institution(kinsen kashitsuke gyō). They were generally used by business owners, and their interest rates were significantly higher than those charged by pawnshops (Shibuya 2000, pp. 184; 248). The Tokyo Institute for Municipal Research expressed concern over the growing number of unscrupulous moneylenders, many of whom were arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp. 106–108).
Credit unions (shinyō kumiai) were another option available to working-class households. In Tokyo City, there were $215$ unions in 1923. However, according to a survey of $184$ of these unions, the number of members from the manufacturing sector was only $5,662$ (Tokyo Institute for Municipal Research 1925c, p. 20). This represented a small share of male workers in the manufacturing sector, as measured in the 1920 Population Census.
@@ A.3 Development of Tsukishima
Figure A.3: Administrative Wards in Tokyo City
Notes: The name in each lattice indicates the administrative ward in Tokyo. The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima. The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River. The borders of each ward are based on the administrative districts in 1920. Geographical coordinate system is based on JGD2000/(B, L). Source: Created by the author using the official shapefile (Ministry of Land, Infrastructure, Transport and Tourism, database).
Figure A.4: Map of Tsukishima: Ishikawajima, Tsukudajima, Shin-tsukudajima, and Lots 1–3
@@
Notes: Tsukishima contains Ishikawajima, Tsukudajima, Shin-tsukudajima, and Tsukishima Lots 1–3. This map does not illustrate the northernmost part of Ishikawajima (see Figure 2 for this part). “Ferries” shows the Tsukishima to the mainland of Kyobashi Ward sea routes called Tsukuda, Tsukishima, and Kachidoki routes (right to left). “Aioi bridge” links Shin tsukudajima and Fukagawa ward. Source: Department of Health, Ministry of the Interior 1923c, 15th map. The author modified and adjusted the tone using Adobe Photoshop 24.7.0.
@@
Tsukishima (translated as “moon island”) is a reclaimed area on a sandbar of the Sumida River (Figure A.3). [^68] The following description in this section is based on Chūō Ward (1994, pp. 122–124). The islands of Ishikawajima and Tsukudajima at the northern end mark the starting point of Tsukishima’s territory (Figure A.4). Ishikawajima was established in 1626 by Hachizaemon Ishikawa, a boatman who received a fief from the shogunate. In 1790, Ishikawajima Prison (ninsoku yoseba) was constructed to accommodate minor offenders and homeless individuals, providing a facility for reintegration into society. After Matthew Perry’s arrival in Japan in 1853, the shogunate instructed the Mito Clan to construct a shipyard on Ishikawajima. A groundbreaking ceremony was held a month later, and a Western-style wooden sailing ship was delivered to the shogunate in 1856. In 1866, the Ishikawajima Beacon was built on the west bank of the prison to signal the position of ships entering the Nihonbashi area. In 1870, the prison was converted into an official place of imprisonment (toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877. The facility was eventually relocated to Sugamo in Kitatoshima County, Tokyo Prefecture, and the penitentiary on Ishikawajima was closed in 1895.
@@
Adjacent to Ishikawajima lies Tsukudajima, where fishermen from Tsukuda in Osaka Prefecture settled in 1644 after receiving land from the shogunate. In 1645, the Tsukuda Ferry began operations. Throughout the Edo period, Tsukudajima became known as a fishing hub, supplying whitefish for the shogun’s household. In 1863, the Tsukudajima Battery was constructed on the southern side of the island as one of $13$ batteries built to defend Edo Bay. This site later became the starting point for the reclamation of Lot 1 of Tsukishima.
@@
In 1884, the Port of Tokyo Waterway Dredging Project began to secure navigable passages by clearing sediment from the shipping channel, and the reclaimed land of Tsukishima was created using the dredged deposits. Lot 1 was completed in 1891, and the Tsukishima Ferry began operations the same year. Lot 2 was reclaimed in 1894, and Shin-tsukudajima (literally, new Tsukudajima) was established in 1896. From the western shore of Shin-tsukudajima, the island offers views of the Bōsō Peninsula in Chiba Prefecture, making it a favored retreat for writers. Cultural figures such as Tōson Shimazaki and Kaoru Osanai stayed there from 1907 through the early Taishō period. In 1903, the completion of the Aioi Bridge improved transportation access and brought water and electricity to Tsukishima, contributing to the growth of the local industry. The Kachidoki Ferry began serving Lot 2 in 1905, and Lot 3 was finally completed in 1913.
@@ A.4 Housing in Tsukishima
Table A.3 summarizes the housing types recorded in the housing survey (Table B.1). As shown, $88$ detached homes (4646%) were located in Tsukudajima. Tsukudajima is reported to be “superior to other areas in terms of housing quality, and the percentage of inferior housing is meager.” [^69] The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp. 69–76). By contrast, Shin-tsukudajima had a low percentage of detached homes. According to the report, it was “inhabited by unskilled laborers” and “occupied the most inconvenient housing status of all the islands.” Lot 1, considered the center of life in Tsukishima, consisted mainly of numerous tenement-style residences. Lot 2, by comparison, had a relatively high percentage of single-family and two-family houses and appears to have been characterized more as a “workers’ residential area” than Lot 1. These findings suggest possible geographic differences in income distribution reflected in housing type. Despite this, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) were row houses, while only $190$ were single-family homes. This indicates that most workers lived in similar tenements and experienced comparable living standards. Figure A.5 shows a photo of these tenement houses.
@@
Note: Figures A.5a and A.5b show the two- and one-story tenements in Tsukishima, respectively. Source: Department of Health, Ministry of the Interior 1923c, figures 62–63; 66.
@@ A.5 Machinery Factories in Tsukishima
First, I use the results of the Primary School Children Survey, included in the Tsukishima Survey, to analyze the occupations of household heads. This survey was conducted among 4th–6th graders at Tsukishima 1st and 2nd Elementary Schools in July 1919. A total of $1,173$ children were in these grades, and the survey covered all $1,100$ children residing in Tsukishima. Since no other elementary schools existed, this represents a complete survey of all 4th–6th graders in Tsukishima. Although it excludes households without children in these grades, it remains a valuable source for understanding households with children.
Panel A of Table A.4 presents the industrial structures. The composition ratios are similar to the census-based statistics shown in Panel A of Table 1. The report does not clarify whether siblings in the same household are counted more than once. Nevertheless, the similarity between the results confirms that this survey provides reliable statistics for identifying overall trends in family households in Tsukishima.
Panel B of Table A.4 classifies the “social classes” of household heads working in the manufacturing sector into six categories. According to the report, large business owners are classified as the capitalist class; small business owners, free enterprise owners, and directors are grouped as the middle class; and workers are defined as the working class (Department of Health, Ministry of the Interior 1923a, p. 50). Most household heads were workers, with only 1% in the capitalist class and less than 20% in the middle class. Although the definition of “executive” is unclear, “small business owners” appear to refer to small factory owners whose earnings are similar to those of skilled worker households (Department of Health, Ministry of the Interior 1923a, p. 64).
@@
Notes: This table summarizes the statistics from the complete survey on the 4th–6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Panel A summarizes the heads’ industries, and Panel B sorts the “worker type” of the heads in the manufacturing sector, respectively. Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
@@
In November 1920, $214$ factories existed in Tsukishima, of which $168$ (approximately 8080%) were classified as machinery factories. The official report notes that wooden pattern factories could be included in the machinery sector because they were part of the casting processes (Department of Health, Ministry of the Interior 1923a, p. 388). Including the $24$ wooden pattern factories raises the share of machinery factories to approximately 9090%. Panel A of Table A.5 uses data from the manufacturing survey in the Tsukishima Survey to classify the $168$ machinery factories by size and industrial sector. First, more than 9090% of the plants were concentrated in three sectors: machinery and equipment manufacturing, can manufacturing, and blacksmithing, with only $15$ plants in the casting sector. Among the machinery factories, $113$ (6767%) employed fewer than $15$ workers. Second, many smithing factories also had fewer than $15$ workers. Hence, almost the entire smithing industry in Tsukishima was classified as small businesses. [^70] The official report indicates that $43$ of the $47$ smithing factories were small factories and workplaces called (machi kōba). The report does not clearly define machi kōba. However, since their number consistently exceeds the count of factories with 15 or fewer workers, they likely include smaller factories operated as small businesses (Department of Health, Ministry of the Interior 1923a, pp. 392; 410–411).
@@
Panel B of Table A.5 classifies $620$ workers based on the primary school survey data used in Table A.4. Of these, $455$ are skilled workers and $165$ are unskilled. Among the unskilled workers, 45 worked in factories and cottage industries, while 120 were day laborers considered low-income workers. According to the same report, $490$ of the $620$ workers were industrial workers. Day laborers are classified as non-industrial workers. This is consistent with the calculation: $620$ workers minus $120$ day laborers equals $490$ industrial workers. Therefore, approximately 9090% of industrial workers ($455/490$) can be regarded as skilled workers: 6565% worked in factories, 1717% in cottage industries, and 1818% were craftsmen. Since workers in cottage and handicraft industries are classified as working in small factories, roughly 6565% of these skilled workers likely worked in medium- to large-scale factories (with $15+$ workers).
@@
Notably, this survey includes only household heads with primary school children aged 9–11 (i.e., 4th to 6th grades). This suggests that most household heads were in their thirties, an age range typically associated with skilled workers. Therefore, the share of skilled workers among the entire working population may be lower than 9090%. Nevertheless, the data confirm that most Tsukishima workers in their 30s were skilled.
@@
Notes: Panel A summarizes the number of machinery factories by the scale and type of the factory. Statistics are from the manufacturing census conducted in the Tsukishima Survey (November 1, 1920). The threshold of the scale (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916. Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th–6th graders in Tsukishima 1st and 2nd Primary Schools in July 1919. Sources: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
@@ A.6 Works in the Machinery Factories
Panel A of Table A.5 summarizes the number of machinery factories by plant type. Machinery and equipment manufacturing factories accounted for 3939% of all machinery factories. Typical jobs in these factories included finishers (or fitters) and turners. A finisher is defined as “a skilled worker responsible for finishing, assembling, and commissioning machine components” and is considered “one of the finest skilled workers because the role requires proficiency in all aspects to correct defects in machine making” (Department of Health, Ministry of the Interior 1923a, p. 393). The turner is described as the “middle axis of machinists,” a category that includes most semi-skilled and skilled workers. Becoming a fully qualified finisher or lathe operator typically took at least five years. Apprentices usually entered the workforce between ages $14$ and $18$, but did not remain in the same factory throughout their apprenticeship. After acquiring basic skills for 1–2 years, they often moved to another factory to further develop their expertise (Department of Health, Ministry of the Interior 1923a, p. 394). Wages were generally paid either on a piece-work or time basis. In Tsukishima, time contracts were widely adopted and payment followed the “residual time” method, whereby “if the work is completed within the contracted time, the remaining time is treated as a bonus.” For example, if $12$ hours are contracted but the work is completed in $10$ hours, the remaining $2$ hours are paid as a bonus (Department of Health, Ministry of the Interior 1923a, pp. 396-397). Wages were distributed twice a month, divided into two payment periods. Average working hours were reported to be approximately $10$ hours per day, with occasional all-night shifts (Department of Health, Ministry of the Interior 1923a, p. 402). Overtime averaged approximately two hours per day (Department of Health, Ministry of the Interior 1923a, p. 402).
@@
Next, I provide a summary of the workforce in the canning factories, which accounted for 2424% of all factories. Canning factories were engaged in shipbuilding, constructing bridges, manufacturing steel frames for construction, and other work, including the fabrication and repair of gas reservoirs (Department of Health, Ministry of the Interior 1923a, p. 406). Workers in these factories were called “boiler makers” and typically bent steel plates and joined them with rivets. They were divided into rivetter, holder-up, plater, angle-ironsmith, and caulker. Work was usually performed in groups, led by skilled angle-ironsmiths and caulkers. The “bōsin” of the rivetter served as the group leader and provided instructions. Wages were paid on a piece-rate basis, with a fixed daily wage during the off-season.
@@
Smithing was the second most common industry, accounting for 2828% of the total. Panel A of Table A.5 shows that most smithing operations were relatively small. Workers produced boulders, nuts, rivets, lathe stops, jigs, and spanners for can manufacturing and construction. These small shops often acted as subcontractors supplying metal parts to larger factories (Department of Health, Ministry of the Interior 1923a, p. 412). Three main jobs existed: yokoza, sakite, and tatara-fuki (bellows-blower). Yokoza referred to self-employed skilled workers (oyakata), while the other two referred to young, unskilled laborers. In small operations, wages were usually determined as a percentage through voluntary agreement with the proprietor, and working hours were irregular (Department of Health, Ministry of the Interior 1923a, pp. 411–413).